# Impact of positrons on electrical conductivity of hot and dense astrophysical plasma

### v1

@@ Abstract
We study the influence of positrons on the outer crusts of neutron stars and the interiors of white dwarfs, introducing them as a novel component in both the composition of matter and in transport processes. We solve a system of coupled Boltzmann kinetic equations for the electron and positron distribution functions in the relaxation-time approximation, taking into account electron–ion, positron–ion, and electron–positron collisions. The relevant scattering matrix elements are calculated from one-plasmon exchange diagrams, with in-medium polarization tensors derived within the hard–thermal–loop effective theory. Numerical results are obtained for matter composed of carbon nuclei. We find that the conductivity rises with temperature, following a power law $\sigma\propto T^{4}$ in the semi-degenerate regime and $\sigma\propto T$ in the nondegenerate regime, due to the intense creation of thermal electron–positron pairs and the resulting collisions among them. These results highlight the importance of including positrons in the transport properties of heated, dense astrophysical plasmas.
@@ I Introduction
Electrical conductivity of matter at subsaturation densities in neutron stars and cores of white dwarfs play a central role in several astrophysical problems, such as the magnetic field evolution 1992ApJ…395..250G ; 1995JApA…16..217B ; 1996ApJ…458..347M ; Ascenzi2024 , dissipation of magnetohydrodynamic waves 1979ApJ…227..995E ; 2014Ap…..57..530S , etc. Conductivity of dense neutron-star matter has been studied extensively in the cold regime relevant to isolated stars ($T\leq 1$ MeV) Flowers1976 ; Yakovlev1980 ; Urpin1980 ; Flowers1981 ; Itoh1983 ; Itoh1984 ; Nandkumar1984MNRAS ; Sedrakian1987 ; Itoh1993 ; Baiko1998 ; Potekhin1999 ; Shternin2006 ; Itoh2008 , see Ref. Schmitt2018 for a recent review. However, formation of hot crustal matter in compact stars is anticipated by various astrophysical scenarios, such as supernova explosions leading to hot protoneutron stars, remnants of binary neutron star mergers, as well as heating of neutron star crusts in accreting binaries. In this regime, the plasma forms a liquid state of correlated ions, while the relativistic electron gas becomes non-degenerate. This transition triggers intense production of electron–positron pairs, introducing a positron component into the plasma that contributes significantly to its conductivity. Finite temperature properties of such plasma are expected to significantly deviate from zero‑temperature (cold, degenerate) results.
@@
The present work extends the analysis of Ref. Harutyunyan2016 to a hot plasma state in which a non-negligible population of positrons coexists with electrons. Although positron contributions were neglected in earlier studies, they become critically important in the outer crust at temperatures $T\gtrsim 1\textrm{MeV}=1.16\times 10^{10}$ K. In this regime, electrons and positrons constitute the primary charge carriers, and electrical conduction is governed mainly by electron–ion, positron–ion, and electron–positron scattering processes.
To describe transport, we employ Boltzmann kinetic theory in the relaxation-time approximation to obtain the non-equilibrium distribution functions of electrons and positrons, fully accounting for electron–ion, positron–ion, and electron–positron collisions. In-medium screening effects are incorporated through the structure factor of the one-component classical plasma and the hard-thermal-loop effective-theory treatment of the plasmon self-energy Braaten1990 ; Braaten1992 ; Harutyunyan2016 .
This work is structured as follows. In Sec. II we derive the electrical conductivity for a multicomponent plasma consisting of electrons, positrons and ions, solving the system of linearized coupled Boltzmann equations for electron and positron distribution functions. In Section III, the matrix elements for electron-ion, positron-ion and positron-electron scattering processes and the resulting collision rates are discussed. Section IV presents the phase structure of crustal matter and numerical results for electrical conductivity. Our results are summarized in Sec. V. Appendix A provides the details of the derivation of the electron-positron scattering matrix element, and Appendix B computes the collision rate due to this scattering channel.
@@ II Electrical conductivity from the Boltzmann equation
$\frac{\partial f^{-}}{\partial t}+\bm{v}\frac{\partial f^{-}}{\partial\bm{r}}-e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{-}}{\partial\bm{p}}=I^{-},$ (1)
$\frac{\partial f^{+}}{\partial t}+\bm{v}\frac{\partial f^{+}}{\partial\bm{r}}+e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{+}}{\partial\bm{p}}=I^{+},$ (2)
where $f^{-}$ and $f^{+}$ are the distribution functions for electrons and positrons, respectively, $\bm{E}$} and $\bm{H}$} are the electric and magnetic fields, $\bm{v}$} is the electron/positron velocity, and $I^{-}$ and $I^{+}$ are the collision integrals, respectively, for electrons and positrons, which include electron-ion and electron-positron scattering for $I^{-}$, and positron-ion and positron-electron scattering for $I^{+}$
@@
where the sums in the first terms go over the type of ions (here we consider the general case of a multi-component plasma, where more than one sort of ions can be present simultaneously). For electron-ion and positron-ion collision integrals, we have
@@
where the indices $1$ and $3$ label the incoming and outgoing electrons/positrons, respectively, $g^{i}_{2,4}\equiv g^{i}(p_{2,4})$ are the ion distribution functions before and after the collision, and ${\cal M}_{12\to 34}^{ei,pi}$ is the electron/positron–ion scattering matrix element. The upper signs in Eqs. (5) and (6) correspond to Bose statistics and the lower signs to Fermi statistics for ions. We also use the shorthand notation $\sum\limits_{i}=\int d{\bm{p}}_{i}/(2\pi)^{3}$. Assuming that ions remain in thermal equilibrium, we take $g^{i}(p)$ in the form of the Bose or Fermi distribution function, $g^{i}(p)=\left[e^{\beta(\varepsilon_{p}-\mu_{i})}\mp 1\right]^{-1},$ where $\varepsilon_{p}=p^{2}/M_{i}$, and $M_{i}$ and $\mu_{i}$ are the mass and chemical potential of ions of type $i$, while $\beta=T^{-1}$ is the inverse temperature.
@@
where $f^{0\pm}({\epsilon})=[e^{\beta(\epsilon\pm\mu_{e})}+1]^{-1}$ are the local Fermi distribution functions for positrons/electrons with $\epsilon=\sqrt{p^{2}+m^{2}}$, $m$ is the electron mass, $\mu$ is the electron chemical potential, $\delta f^{\pm}\ll f^{0\pm}$ are small perturbations, and $\phi^{\pm}$ are unknown functions; additionally, introducing fermionic velocity $\bm{v}=\partial\epsilon/\partial\bm{p}=\bm{p}/\epsilon$, we quote some of the derivatives that will be used below
$\frac{\partial f^{0\pm}}{\partial\bm{p}}$ $=$ $\bm{v}\frac{\partial f^{0\pm}}{\partial\epsilon},\quad\frac{\partial f^{0\pm}}{\partial\epsilon}=-\beta f^{0\pm}(1-f^{0\pm}).$ (10)
We assume plasma is spatially uniform, time-independent and external fields are uniform and constant, i.e., there is no spatial variation of the fermionic distribution function, with only momentum space evolution.
We next linearize the Boltzmann equation by substituting Eq. (9) into Eqs. (1) and (2) and keeping only the terms linear in $\delta f^{\pm}$ and in the electric field $\bm{E}$}, which implies that in the terms proportional to $\bm{E}$}, the substitutions $f^{\pm}\to f^{0\pm}$ can be made. We further neglect magnetic fields, which induce anisotropy in the conductivity; for the discussion of the effects of $\bm{H}$} field on transport, see Ref. Harutyunyan2016 . Thus, the linearized Boltzmann equations take the form
@@
We seek solutions of Eqs. (11) and (12) in the form
@@
and Eqs. (11) and (12) become
@@
$\hskip-34.14322pt\times\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big),\quad$
@@
Similar expressions hold for positons with the replacements $ei\to pi$, $ep\to pe$ and $f^{0-}\to f^{0+}$. These quantities represent appropriately weighted mean collision rates between the corresponding particles. Solving (29) and (30) then yields
@@
where the collision rates on different types of nuclei sum up.
In the regimes where electron-positron collisions become important, the number of electrons and positrons is almost equal, and we find that practically always $\nu_{ep}\simeq\nu_{pe}$, such that $\nu_{ep}-\nu_{pe}\ll\nu_{ei}\simeq\nu_{pi}$. Then Eqs. (33) and (34) can be simplified to
$X^{-}$ $=$ $\frac{\sum_{i}\nu_{pi}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})},$ (36)
$X^{+}$ $=$ $\frac{\sum_{i}\nu_{ei}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})}.$ (37)
In the opposite limit of fast electron-positron scattering we have $\nu_{ei}\simeq\nu_{pi}\ll\nu_{ep}\simeq\nu_{pe}$, therefore
@@
which can be derived also directly from Eqs. (29) and (30).
We next write explicit expressions for the mean collision rates by substituting Eqs. (13)–(16) into (31) and (32) (and their analogues for positrons) to obtain
@@
where we used Eqs. (9) and (17), and defined the conductivity matrices (the factor 2 accounts for spin degeneracy) as
@@
with total conductivity ${\sigma}={\sigma}^{+}+{\sigma}^{-}$. The electrical conductivity is thus fully determined once the relaxation times $\tau^{\pm}$ are known. Employing the ansatz $\tau^{\pm}=X^{\pm}\epsilon$, Eq. (47) becomes
@@ III Scattering amplitudes and collision rates
III.1 The electron-positron scattering matrix element
@@
$\tilde{J}^{\mu}=-e^{*}\bar{v}_{2}\gamma^{\mu}u_{1},\quad\tilde{J}^{\prime\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}v_{4},$ (56)
where $e^{*}=\sqrt{4\pi}e$, $(\omega,{\bm{q}})=p_{1}-p_{3}=p_{4}-p_{2}$, $(\tilde{\omega},\tilde{{\bm{q}}})=p_{1}+p_{2}=p_{3}+p_{4}$, $u_{a}={u}^{s_{a}}(p_{a})$, $v_{b}={v}^{s_{b}}(p_{b})$, $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of these currents transverse to $\bm{q}$, and $\tilde{\bm{J}}_{\perp},\tilde{\bm{J}}^{\prime}_{\perp}$ are those transverse to $\tilde{\bm{q}}$. The screening of the interaction is taken into account in terms of the longitudinal $\Pi_{L}(\omega,q)$ and transverse $\Pi_{T}(\omega,q)$ components of the polarization tensor.
Details of the squaring and spin-averaging of the $ep$ matrix element are given in Appendix A. In this work, we neglect the interference terms between the $t$- and $s$-channels in the electron–positron scattering amplitude. The final results for the $t$- and $s$-channel spin-averaged scattering amplitudes are
@@
where using the shorthands $\cos\alpha=x_{\alpha}$ and $\cos\theta=x_{\theta}$,
@@
III.2 The electron-ion and positron-ion scattering matrix elements
In analogy to Eq. (50), the one-plasmon-exchange scattering amplitudes of electrons/positrons off ions can be written as (see Fig. 2)

### v2

@@ Abstract
We study the influence of positrons on the outer crusts of neutron stars and the interiors of white dwarfs, introducing them as a novel component in both the composition of matter and in transport processes. We solve a system of coupled Boltzmann kinetic equations for the electron and positron distribution functions in the relaxation-time approximation, taking into account electron–ion, positron–ion, and electron–positron collisions. The relevant scattering matrix elements are calculated from one-plasmon exchange diagrams, with in-medium polarization tensors derived within the hard–thermal–loop effective theory. Numerical results are obtained for matter composed of carbon, iron and helium nuclei. We find that the conductivity rises with temperature, following a power law $\sigma\propto T^{4}$ in the semidegenerate regime and $\sigma\propto T$ in the nondegenerate regime, due to the intense creation of thermal electron–positron pairs and the resulting collisions among them. These results highlight the importance of including positrons in the transport properties of heated, dense astrophysical plasmas.
@@ I Introduction
Electrical conductivity of matter at subsaturation densities in neutron stars and cores of white dwarfs plays a central role in several astrophysical problems, such as the magnetic field evolution 1992ApJ…395..250G ; 1995JApA…16..217B ; 1996ApJ…458..347M ; Ascenzi2024 , dissipation of magnetohydrodynamic waves 1979ApJ…227..995E ; 2014Ap…..57..530S , etc. The conductivity of dense neutron-star matter has been studied extensively in the cold regime relevant to isolated stars ($T\leq 1$ MeV) Flowers1976 ; Yakovlev1980 ; Urpin1980 ; Flowers1981 ; Itoh1983 ; Itoh1984 ; Nandkumar1984MNRAS ; Sedrakian1987 ; Itoh1993 ; Baiko1998 ; Potekhin1999 ; Shternin2006 ; Itoh2008 ; see Ref. Schmitt2018 for a recent review. However, formation of hot crustal matter in compact stars is anticipated by various astrophysical scenarios, such as supernova explosions leading to hot protoneutron stars, remnants of binary neutron star mergers, as well as heating of neutron star crusts in accreting binaries. In this regime, the plasma forms a liquid state of correlated ions, while the relativistic electron gas becomes nondegenerate. This transition triggers intense production of electron–positron pairs, introducing a positron component into the plasma that contributes significantly to its conductivity. Finite temperature properties of such a plasma are expected to significantly deviate from zero‑temperature (cold, degenerate) results.
@@
The present work extends the analysis of Ref. Harutyunyan2016 to a hot plasma state in which a non-negligible population of positrons coexists with electrons. Although positron contributions were neglected in earlier studies, they become critically important in the outer crust of neutron stars at temperatures $T\gtrsim 1\textrm{MeV}=1.16\times 10^{10}$ K. In this regime, electrons and positrons constitute the primary charge carriers, and electrical conduction is governed mainly by electron–ion, positron–ion, and electron–positron scattering processes.
To describe transport, we employ Boltzmann kinetic theory in the relaxation-time approximation to obtain the nonequilibrium distribution functions of electrons and positrons, fully accounting for electron–ion, positron–ion, and electron–positron collisions. In-medium screening effects are incorporated through the structure factor of the one-component classical plasma and the hard-thermal-loop effective-theory treatment of the plasmon self-energy Braaten1990 ; Braaten1992 ; Harutyunyan2016 .
This work is structured as follows. In Sec. II we derive the electrical conductivity for a multicomponent plasma consisting of electrons, positrons and ions, solving the system of linearized coupled Boltzmann equations for electron and positron distribution functions. In Sec. III, the matrix elements for electron-ion, positron-ion, and positron-electron scattering processes and the resulting collision rates are discussed. Section IV presents the phase structure of crustal matter and numerical results for electrical conductivity. Our results are summarized in Sec. V. Appendix A provides the details of the derivation of the electron-positron scattering matrix element, and Appendix B computes the collision rate due to this scattering channel.
@@ II Electrical conductivity from the Boltzmann equation
$\frac{\partial f^{-}}{\partial t}+\bm{v}\frac{\partial f^{-}}{\partial\bm{r}}-e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{-}}{\partial\bm{p}}=I^{-},\quad$ (1)
$\frac{\partial f^{+}}{\partial t}+\bm{v}\frac{\partial f^{+}}{\partial\bm{r}}+e(\bm{E}+\bm{v}\times\bm{H})\frac{\partial f^{+}}{\partial\bm{p}}=I^{+},\quad$ (2)
where $f^{-}$ and $f^{+}$ are the distribution functions for electrons and positrons, respectively, $\bm{E}$} and $\bm{H}$} are the electric and magnetic fields, $\bm{v}$} is the electron or positron velocity, and $I^{-}$ and $I^{+}$ are the collision integrals, respectively, for electrons and positrons, which include electron-ion and electron-positron scattering for $I^{-}$, and positron-ion and positron-electron scattering for $I^{+}$,
@@
where the sums in the first terms run over the ion species, allowing for the general case of a multicomponent plasma in which several types of ions may coexist. In general, the composition is described by a statistical distribution of nuclear species; however, in the present work, we assume that a single species is dominant, and we restrict our numerical computations to a plasma composed of this primary nuclear species. This approximation is valid in the low- to intermediate-temperature regime considered here, but may require corrections at higher temperatures, where differences in the binding energies of various nuclei become less significant in determining the composition of matter. A detailed study of a genuinely multicomponent plasma with temperature-dependent compositions is left for future work.
For electron-ion and positron-ion collision integrals, we have
@@
where the indices 11 and $3$ label the incoming and outgoing electrons or positrons, respectively, $g^{i}_{2,4}\equiv g^{i}(p_{2,4})$ are the ion distribution functions before and after the collision, and ${\cal M}_{12\to 34}^{ei,pi}$ is the electron- or positron–ion scattering matrix element. The upper signs in Eqs. (5) and (6) correspond to Bose statistics and the lower signs to Fermi statistics for ions. We also use the shorthand notation $\sum\limits_{i}=\int d{\bm{p}}_{i}/(2\pi)^{3}$. Assuming that ions remain in thermal equilibrium, we take $g^{i}(p)$ in the form of the Bose or Fermi distribution function, $g^{i}(p)=\left[e^{\beta(\varepsilon_{p}-\mu_{i})}\mp 1\right]^{-1},$ where $\varepsilon_{p}=p^{2}/M_{i}$, and $M_{i}$ and $\mu_{i}$ are the mass and chemical potential of ions of type $i$, while $\beta=T^{-1}$ is the inverse temperature.
@@
where $f^{0\pm}({\epsilon})=[e^{\beta(\epsilon\pm\mu_{e})}+1]^{-1}$ are the local Fermi distribution functions for positrons $(+)$ and electrons ()(-) with $\epsilon=\sqrt{p^{2}+m^{2}}$, $m$ is the electron mass, $\mu_{e}$ is the electron chemical potential, $\delta f^{\pm}\ll f^{0\pm}$ are small perturbations, and $\phi^{\pm}$ are unknown functions; additionally, introducing fermionic velocity $\bm{v}=\partial\epsilon/\partial\bm{p}=\bm{p}/\epsilon$, we give some of the derivatives that will be used below,
$\frac{\partial f^{0\pm}}{\partial\bm{p}}$ $=$ $\bm{v}\frac{\partial f^{0\pm}}{\partial\epsilon},\quad\frac{\partial f^{0\pm}}{\partial\epsilon}=-\beta f^{0\pm}(1-f^{0\pm}).\quad$ (10)
We assume plasma is spatially uniform, time independent and external fields are uniform and constant; i.e., there is no spatial variation of the fermionic distribution function, with only momentum space evolution.
We next linearize the Boltzmann equation by substituting Eq. (9) into Eqs. (1) and (2) and keeping only the terms linear in $\delta f^{\pm}$ and in the electric field $\bm{E}$}, which implies that in the terms proportional to $\bm{E}$}, the substitutions $f^{\pm}\to f^{0\pm}$ can be made. We further neglect magnetic fields, which induce anisotropy in the conductivity; for the discussion of the effects of the $\bm{H}$} field on transport, see Ref. Harutyunyan2016 . Thus, the linearized Boltzmann equations take the form
@@
We search for the solutions of Eqs. (11) and (12) in the form
@@
and Eqs. (11) and (12) become (the cancellation of the vector $\bm{E}$ from both sides of the equations is justified as these should be satisfied for an arbitrary direction of $\bm{E}$; equivalently, one may project the equation along $\bm{E}$ and divide by $|\bm{E}|$)
@@
$\hskip-34.14322pt\times\Big(\tau^{+}_{1}\bm{v}_{1}-\tau^{-}_{2}\bm{v}_{2}-\tau^{+}_{3}\bm{v}_{3}+\tau^{-}_{4}\bm{v}_{4}\Big),\qquad$
@@
Similar expressions hold for positrons with the replacements $ei\to pi$, $ep\to pe$ and $f^{0-}\to f^{0+}$. These quantities represent appropriately weighted mean collision rates between the corresponding particles. Solving Eqs. (29) and (30) then yields
@@
where the collision rates on different types of nuclei are summed up.
In the regimes where electron-positron collisions become important, the numbers of electrons and positrons are almost equal, and we find that, practically always, $\nu_{ep}\simeq\nu_{pe}$, such that $\nu_{ep}-\nu_{pe}\ll\nu_{ei}\simeq\nu_{pi}$. Then, Eqs. (33) and (34) can be simplified to
$X^{-}$ $=$ $\frac{\sum_{i}\nu_{pi}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})},\quad$ (36)
$X^{+}$ $=$ $\frac{\sum_{i}\nu_{ei}}{\big(\sum_{i}\nu_{ei}\big)\big(\sum_{i}\nu_{pi}\big)+\nu_{ep}\sum_{i}(\nu_{ei}+\nu_{pi})}.\quad$ (37)
In the limit of fast electron-positron scattering we have also $\nu_{ei}\simeq\nu_{pi}\ll\nu_{ep}\simeq\nu_{pe}$; therefore,
@@
which can also be derived directly from Eqs. (29) and (30).
We next write explicit expressions for the mean collision rates by substituting Eqs. (13)–(16) into (31) and (32) (and their analogs for positrons) to obtain
@@
where we used Eqs. (9) and (17) and defined the conductivity matrices (the factor 2 accounts for spin degeneracy) as
@@
with total conductivity ${\sigma}={\sigma}^{+}+{\sigma}^{-}$. The electrical conductivity is thus fully determined once the relaxation times $\tau^{\pm}$ are known. By employing the ansatz $\tau^{\pm}=\epsilon X^{\pm}$, Eq. (47) becomes
@@
We remark that the range of validity of the Boltzmann kinetic theory employed here can be derived from the common considerations Lifshitz_Pitaevskii1981 . The main assumption is that the collisions are instantaneous, which is justified if the average scattering length which has the order of the Debye screening radius $r_{D}=q_{D}^{-1}$ of the electric field (see Sec. III.2 for details) is much smaller than the electron or positron mean free paths $l^{\pm}\simeq\tau^{\pm}$ between collisions (assuming ultra-relativistic particles with average velocities $\bar{v}^{\pm}\simeq 1$). The particle mean collision times and related mean free paths are supposed to be much smaller than the characteristic length scale $L$ and the inverse frequency $\Omega^{-1}$ of macroscopic perturbations of the electric field. Thus, the general validity condition can be written as
$q_{D}^{-1}\ll\tau^{\pm}\ll{\rm min}\{L,\Omega^{-1}\}.$
Our numerical results (see Sec. IV, Figs. 46) show that in the regime of interest $0.03\leq q_{D}\leq 2$ MeV and $10^{2}\leq\tau^{\pm}\leq 10^{4}$ MeV-1 (recall that 1 s= $1.52\times 10^{21}$ MeV-1); therefore the condition $q_{D}\tau^{\pm}\gg 1$ is satisfied practically in the whole regime of interest with high accuracy.
@@ III Scattering amplitudes and collision rates
III.1 Electron-positron scattering matrix element
@@
$\tilde{J}^{\mu}=-e^{*}\bar{v}_{2}\gamma^{\mu}u_{1},\quad\tilde{J}^{\prime\mu}=-e^{*}\bar{u}_{3}\gamma^{\mu}v_{4}.$ (56)
Here, $e^{*}=\sqrt{4\pi}e$, $(\omega,{\bm{q}})=p_{1}-p_{3}=p_{4}-p_{2}$, $(\tilde{\omega},\tilde{{\bm{q}}})=p_{1}+p_{2}=p_{3}+p_{4}$, $u_{a}={u}^{s_{a}}(p_{a})$, $v_{b}={v}^{s_{b}}(p_{b})$, $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of these currents transverse to $\bm{q}$, and $\tilde{\bm{J}}_{\perp},\tilde{\bm{J}}^{\prime}_{\perp}$ are those transverse to $\tilde{\bm{q}}$. The screening of the interaction is taken into account in terms of the longitudinal $\Pi_{L}(\omega,q)$ and transverse $\Pi_{T}(\omega,q)$ components of the polarization tensor.
Details of the squaring and spin averaging of the $ep$ matrix element are given in Appendix A. In this work, for simplicity, we neglect the interference terms between the $t$- and $s$-channels in the electron–positron scattering amplitude. The final results for the $t$- and $s$-channel spin-averaged scattering amplitudes are
@@
where, using the shorthands $\cos\alpha=x_{\alpha}$ and $\cos\theta=x_{\theta}$,
@@
III.2 Electron-ion and positron-ion scattering matrix elements
In analogy to Eq. (50), the one-plasmon-exchange scattering amplitudes of electrons or positrons off ions can be written as (see Fig. 2)